# FlashVault: Versatile In-NAND Self-Encryption with Zero Area Overhead

### v1

@@ VI-A Block Cipher Engine (BCE)
4) Shift Unit (SU): Some block ciphers use shift operations to increase the complexity of deducing key values from encrypted data. For example, in AES, each row of a 44 data block is shifted by a different offset to shuffle data locations [13]. The SU supports such operations and consists of two barrel shifters. These shifters can be combined to support up to 64-bit arithmetic, logical, and rotate shift operations.
@@ VI-B Asymmetric Cryptographic Engine (ACE)
1) Hash ALU Cluster: Digital signature algorithms typically hash messages, then sign the resulting digest [22]. Hashing enables efficient handling of arbitrary-length messages by reducing them to fixed-size digests. Furthermore, certain PQC schemes like SPHINCS+ [8] rely on hash functions as core primitives. To support conventional signature hashing and hash-based PQC, ACE integrates a dedicated Hash ALU cluster. Hash functions are inherently parallelizable [6] as they process input by dividing it into fixed-size blocks and operating independently via tree hashing [6]. To support this parallelism, FlashVault incorporates eight Hash ALUs to accelerate the hashing process.
Fig. 11-(b) shows the block diagram of the Hash ALU, which comprises four key components optimized for the primitive operations used in cryptographic hash functions: (i) a modular addition unit, (ii) a logic operation unit, (iii) a permutation unit, and (iv) a shift unit. The modular addition unit supports both standard and modular additions for hash state updates. The logic operation unit supports bitwise operations such as XOR, AND, and NOT to introduce non-linearity by disrupting linear relationships between input and output. The permutation unit, based on a Benes network, reorders bit positions to enhance diffusion. The shift unit supports word-level (e.g., 32- or 64-bit) logical and circular shifts, as commonly used in hash functions like SHA-2 [76].
2) Asymmetric Cipher ALUs: PKC and PQC computations typically proceed sequentially, with each stage depending on prior results, e.g., modular exponentiation, scalar multiplication, and NTT/FFT-based transforms. To efficiently support this, ACE integrates two asymmetric cipher ALUs, which include the arithmetic, logic, permutation, and shift units. These units handle high-precision integer arithmetic, bit manipulations, and data transformations. While architecturally similar to the Hash ALU, asymmetric cipher ALUs support more diverse and complex arithmetic for PKC/PQC. That is, the arithmetic unit includes extended functional blocks beyond the modular addition unit of the Hash ALU, including adders, multipliers, and modular operators. The arithmetic unit adopts a fixed 64-bit integer datapath, zero-extending narrower operands for alignment. For high-precision computations, operands are divided into 64-bit limbs (subwords), with limb-wise computation involving carry propagation and cross-limb alignment. The adder in the arithmetic unit is designed with a carry-save structure to efficiently support multi-limb accumulation. To reduce propagation delay, the multiplier uses a Wallace tree [120] and the modular unit uses Barrett reduction [54] with shifters and fixed-coefficient multipliers.
@@ VIII-A Area and Power Analysis of FlashVault
1) Methodology: To evaluate the effectiveness of FlashVault, we synthesized all components shown in Fig. 9 and performed place-and-route (P&R) using Synopsys EDA tools at fast/fast and slow/slow corners in a 65nm CMOS process. For low-power design, we utilized the unified power format (UPF) to implement power gating. Additionally, for dynamic power minimization, we implemented clock gating at the RTL level. On-chip memories were instantiated using commercial memory IPs. Area was extracted using Synopsys IC Compiler II [109] and consumed power was estimated using PrimeTime PX [110] based on switching activity interchange format (SAIF) generated from post-layout simulations with StarRC-extracted parasitics [111]. To report results under more advanced technology assumptions, we scaled the area and power consumption to 14nm technology node using scaling factors from [40]. This node selection reflects the technology used in recent commercial SSDs, which commonly adopt nodes in the mid-10nm range, i.e., 14nm [61, 31, 25].
@@ VIII-B Experimental Setup for Architectural Evaluation
1) FlashVault Implementation: As described in Section VIII-A, FlashVault integrates two cryptographic engines within the area and power budgets of Samsung’s 7th-generation V-NAND. Based on this configuration, we evaluate the in-die cryptographic performance. FlashVault is implemented using SimpleSSD [32], incorporating the P&R results in Table II and the system parameters in Table III.
2) Baseline Systems: We compare FlashVault with two baselines: (i) a Near-Core Processing (NCP) architecture and (ii) a real-machine system (CPU + SSD). The NCP architecture uses the same cryptographic engines as FlashVault with identical configuration, but places them near the SSD controller instead of inside the NAND die. Following recent NCP designs [49, 62, 85, 66], our NCP baseline uses a device DRAM as working memory for encryption. The real-machine system performs encryption on the host CPU which follows the hardware setup provided in Section III-B1.
3) Methodology:We modified the hardware abstraction layer of SimpleSSD [32] to enable end-to-end latency estimation for NVMe requests. To reflect the latency overhead of cryptographic operations in FlashVault and NCP, we injected cycle-level latency estimated from design-for-test (DFT) verification using Synopsys tools into the I/O request handling path and controller-side I/O processing path of SimpleSSD, respectively. In addition, we modified the DRAM abstraction layer in the NCP model to reflect the DRAM access latency during cryptographic operations. The real-machine system follows the methodology described in Section III-B1, with SSD latency derived from baseline SimpleSSD (without cryptographic engines) for fair comparison. Our evaluation uses the block ciphers, PKC, and PQC algorithms listed in Section III-B as benchmarks, assessed by varying the input sizes.
@@ VIII-C Performance Evaluation
3) Signature Generation: Fig. 14 shows the signature generation latency for 1 KB and 10 KB inputs. FlashVault achieves 2.9 and 3.1 lower latency on average than the real-machine system for 1 KB and 10 KB, respectively. Anti-tamper logging, a representative application of signature generation, is typically required to complete within 2 to 15 ms in commercial SSDs, as discussed in Section III-B4. Specifically, real-time embedded systems typically require completion within 2 ms [99], while general-purpose systems tolerate up to 15 ms [1]. However, the real-machine system exceeds the latency bounds of both systems when executing PQC algorithms. In contrast, FlashVault meets the general-purpose latency bounds for all PQC algorithms and satisfies the stricter embedded constraint for all PQC algorithms except SPHINCS+. This indicates that FlashVault supports most cryptographic algorithms in both embedded and general-purpose environments. Since the evaluation uses small input sizes representative of typical log entries, the DRAM access overhead in the NCP is negligible, resulting in a modest performance gain of 0.2 to 1.0% for FlashVault over NCP.
@@
4) FTL Overhead: All prior evaluations were conducted under best-case conditions, with the system freshly initialized and no background FTL activities such as GC and WL. In this state, the baseline FTL latency is approximately 4.6 s for a 1 KB input and increases sub-linearly with larger input sizes. This trend is attributed to the proportional growth in processing time for address translation and data buffering. To assess the impact of GC and WL, we implemented a steady-state environment in SimpleSSD [32] by filling the SSD with dummy data and repeatedly overwriting random locations. Table IV presents latency and best-case-relative overhead under steady-state conditions, using the same inputs and workloads as in the previous evaluation. For block ciphers, program operations incur an average FTL overhead of 46.1%, while read operations show a higher 79.3% overhead due to the relatively short NAND read latency, which makes the FTL more dominant in total latency. For public-key operations, signature verification and signing show average overheads of 3.6% and 127.6%, respectively. The verification incurs minimal overhead, as cryptographic latency dominates, whereas signing suffers from high overhead due to shorter execution time and smaller inputs. SPHINCS+ deviates from the trend in signing overheads due to substantial cryptographic latency, which is caused by hash tree traversal and high computational complexity. Although not shown in the table, verification with a 10 MB input and signing with a 1 KB input result in average GC/WL overheads of 5.6% and 87.6%, respectively.

### v2

@@ VI-A Block Cipher Engine (BCE)
4) Shift Unit (SU): Some block ciphers use shift operations to increase the complexity of deducing key values from encrypted data. For example, in AES, each row of a 44 data block is shifted by a different offset to shuffle data locations [aes_shift]. The SU supports such operations and consists of two barrel shifters. These shifters can be combined to support up to 64-bit arithmetic, logical, and rotate shift operations.
@@ VI-B Asymmetric Cryptographic Engine (ACE)
1) Hash ALU Cluster: Digital signature algorithms typically hash messages, then sign the resulting digest [nist_hash]. Hashing enables efficient handling of arbitrary-length messages by reducing them to fixed-size digests. Furthermore, certain PQC schemes like SPHINCS+ [sphincs] rely on hash functions as core primitives. To support conventional signature hashing and hash-based PQC, ACE integrates a dedicated Hash ALU cluster. Hash functions are inherently parallelizable [6] as they process input by dividing it into fixed-size blocks and operating independently via tree hashing [hash_parallel_survey]. To support this parallelism, FlashVault incorporates eight Hash ALUs to accelerate the hashing process.
Fig. 11-(b) shows the block diagram of the Hash ALU, which comprises four key components optimized for the primitive operations used in cryptographic hash functions: (i) a modular addition unit, (ii) a logic operation unit, (iii) a permutation unit, and (iv) a shift unit. The modular addition unit supports both standard and modular additions for hash state updates. The logic operation unit supports bitwise operations such as XOR, AND, and NOT to introduce non-linearity by disrupting linear relationships between input and output. The permutation unit, based on a Benes network, reorders bit positions to enhance diffusion. The shift unit supports word-level (e.g., 32- or 64-bit) logical and circular shifts, as commonly used in hash functions like SHA-2 [sha].
2) Asymmetric Cipher ALUs: PKC and PQC computations typically proceed sequentially, with each stage depending on prior results, e.g., modular exponentiation, scalar multiplication, and NTT/FFT-based transforms. To efficiently support this, ACE integrates two asymmetric cipher ALUs, which include the arithmetic, logic, permutation, and shift units. These units handle high-precision integer arithmetic, bit manipulations, and data transformations. While architecturally similar to the Hash ALU, asymmetric cipher ALUs support more diverse and complex arithmetic for PKC/PQC. That is, the arithmetic unit includes extended functional blocks beyond the modular addition unit of the Hash ALU, including adders, multipliers, and modular operators. The arithmetic unit adopts a fixed 64-bit integer datapath, zero-extending narrower operands for alignment. For high-precision computations, operands are divided into 64-bit limbs (subwords), with limb-wise computation involving carry propagation and cross-limb alignment. The adder in the arithmetic unit is designed with a carry-save structure to efficiently support multi-limb accumulation. To reduce propagation delay, the multiplier uses a Wallace tree [wallace_multiplier] and the modular unit uses Barrett reduction [barrett_reduction] with shifters and fixed-coefficient multipliers.
@@ VIII-A Area and Power Analysis of FlashVault
1) Methodology: To evaluate the effectiveness of FlashVault, we synthesized all components shown in Fig. 9 and performed place-and-route (P&R) using Synopsys EDA tools at fast/fast and slow/slow corners in a 65nm CMOS process. For low-power design, we utilized the unified power format (UPF) to implement power gating. Additionally, for dynamic power minimization, we implemented clock gating at the RTL level. On-chip memories were instantiated using commercial memory IPs. Area was extracted using Synopsys IC Compiler II [synopsys_icc] and consumed power was estimated using PrimeTime PX [synopsys_pt] based on switching activity interchange format (SAIF) generated from post-layout simulations with StarRC-extracted parasitics [rc_synopsys]. To report results under more advanced technology assumptions, we scaled the area and power consumption to 14nm technology node using scaling factors from [scale_factor]. This node selection reflects the technology used in recent commercial SSDs, which commonly adopt nodes in the mid-10nm range, i.e., 14nm [14nm_nand, 4d_nand_survey, nand_roadmap].
@@ VIII-B Experimental Setup for Architectural Evaluation
1) FlashVault Implementation: As described in Section VIII-A, FlashVault integrates two cryptographic engines within the area and power budgets of Samsung’s 7th-generation V-NAND. Based on this configuration, we evaluate the in-die cryptographic performance. FlashVault is implemented using SimpleSSD [modern_ssd_amber], incorporating the P&R results in Table II and the system parameters in Table III.
2) Baseline Systems: We compare FlashVault with two baselines: (i) a Near-Core Processing (NCP) architecture and (ii) a real-machine system (CPU + SSD). The NCP architecture uses the same cryptographic engines as FlashVault with identical configuration, but places them near the SSD controller instead of inside the NAND die. Following recent NCP designs [optimstore, ecssd, flagger, hardi_ssd_isca], our NCP baseline uses a device DRAM as working memory for encryption. The real-machine system performs encryption on the host CPU which follows the hardware setup provided in Section III-B1.
3) Methodology:We modified the hardware abstraction layer of SimpleSSD [modern_ssd_amber] to enable end-to-end latency estimation for NVMe requests. To reflect the latency overhead of cryptographic operations in FlashVault and NCP, we injected cycle-level latency estimated from design-for-test (DFT) verification using Synopsys tools into the I/O request handling path and controller-side I/O processing path of SimpleSSD, respectively. In addition, we modified the DRAM abstraction layer in the NCP model to reflect the DRAM access latency during cryptographic operations. The real-machine system follows the methodology described in Section III-B1, with SSD latency derived from baseline SimpleSSD (without cryptographic engines) for fair comparison. Our evaluation uses the block ciphers, PKC, and PQC algorithms listed in Section III-B as benchmarks, assessed by varying the input sizes.
@@ VIII-C Performance Evaluation
3) Signature Generation: Fig. 14 shows the signature generation latency for 1 KB and 10 KB inputs. FlashVault achieves 2.9 and 3.1 lower latency on average than the real-machine system for 1 KB and 10 KB, respectively. Anti-tamper logging, a representative application of signature generation, is typically required to complete within 2 to 15 ms in commercial SSDs, as discussed in Section III-B4. Specifically, real-time embedded systems typically require completion within 2 ms [emlog_logging], while general-purpose systems tolerate up to 15 ms [tamper_logging_size]. However, the real-machine system exceeds the latency bounds of both systems when executing PQC algorithms. In contrast, FlashVault meets the general-purpose latency bounds for all PQC algorithms and satisfies the stricter embedded constraint for all PQC algorithms except SPHINCS+. This indicates that FlashVault supports most cryptographic algorithms in both embedded and general-purpose environments. Since the evaluation uses small input sizes representative of typical log entries, the DRAM access overhead in the NCP is negligible, resulting in a modest performance gain of 0.2 to 1.0% for FlashVault over NCP.
@@
4) FTL Overhead: All prior evaluations were conducted under best-case conditions, with the system freshly initialized and no background FTL activities such as GC and WL. In this state, the baseline FTL latency is approximately 4.6 s for a 1 KB input and increases sub-linearly with larger input sizes. This trend is attributed to the proportional growth in processing time for address translation and data buffering. To assess the impact of GC and WL, we implemented a steady-state environment in SimpleSSD [modern_ssd_amber] by filling the SSD with dummy data and repeatedly overwriting random locations. Table IV presents latency and best-case-relative overhead under steady-state conditions, using the same inputs and workloads as in the previous evaluation. For block ciphers, program operations incur an average FTL overhead of 46.1%, while read operations show a higher 79.3% overhead due to the relatively short NAND read latency, which makes the FTL more dominant in total latency. For public-key operations, signature verification and signing show average overheads of 3.6% and 127.6%, respectively. The verification incurs minimal overhead, as cryptographic latency dominates, whereas signing suffers from high overhead due to shorter execution time and smaller inputs. SPHINCS+ deviates from the trend in signing overheads due to substantial cryptographic latency, which is caused by hash tree traversal and high computational complexity. Although not shown in the table, verification with a 10 MB input and signing with a 1 KB input result in average GC/WL overheads of 5.6% and 87.6%, respectively.